OpenStudy (anonymous):

HELP PLEASE!!!!!!!!!!!!!!! FAN &&ND MEDAL FORBEST ANSWER!! Wyatt's dog is tethered to a post in his yard. If the rope tied to the dog is 12 feet long, how many square yards does the dog have access to? (Use 3.14 for π.) a)452.16 yd2 b)50.24 yd2 c)75.36 yd2 d)25.12 yd2

OpenStudy (anonymous):

make me a sammich

OpenStudy (lazyboy):

A

OpenStudy (anonymous):

peanut butter and jelly or ham and cheese??? (=

OpenStudy (anonymous):

wait how is it a???

OpenStudy (anonymous):

its A

OpenStudy (lazyboy):

1.A=pi r² A=3.14*144 A=452.1 answer is a. 452.16 yd²

OpenStudy (anonymous):

no its ham sammich

OpenStudy (anonymous):

THANK YOUUUUU!!!

OpenStudy (anonymous):

1.A=pi r² A=3.14*144 A=452.1 answer is a. 452.16 yd² 2 A=1/2*(15+8)h 46=1/2*(23)*h /2 92=23h /23 h=4 answer is a.

OpenStudy (anonymous):

shes wack its not a its ham sammich

OpenStudy (anonymous):

O.O

OpenStudy (anonymous):

wahhh?? I became a fan of magicaltaylorf202 andLazyBoy

OpenStudy (anonymous):

yeah pretty much

OpenStudy (anonymous):

wait u said fan and medal

OpenStudy (lazyboy):

He can only give me a medal

OpenStudy (anonymous):

i know i already gave the medal to lazy boy cause he gave me the answer first.. sorry.. and IMMA GIRL!!

OpenStudy (lazyboy):

oh ok sry :D

OpenStudy (anonymous):

me too!!!!!!

OpenStudy (lazyboy):

there u go

OpenStudy (anonymous):

lol its okayy(:

OpenStudy (anonymous):

thx alot lazyboy

OpenStudy (anonymous):

if i have already given one could i give another one???

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

u can give one to me but not one to lazyboy

OpenStudy (anonymous):

i might put another one up and then i will give you the medal if you can answer it.

OpenStudy (lazyboy):

link me then

OpenStudy (anonymous):

k

OpenStudy (anonymous):

an octagon has equal sides that measure 11 centimeters. its perimeter is 66 centimeters. true false